OpenStudy (aurorab):

solve for x: 2/3 (x-4)= 2x

OpenStudy (aurorab):

@musa19965 help plzz

OpenStudy (aurorab):

btw the 2/3 is supposed to be a fraction

OpenStudy (vjordillo):

Is this the form? \[\frac{ 2 }{ 3 }(x-4)=2x\]

OpenStudy (compassionate):

2/3x - 2.6 = 2x Add the 2.6 Multiply the 2/3

OpenStudy (anonymous):

its easy just multiply both sides by 3 what we get?? we get 2(x-4)=6x now multiply by 2 we get 2x-8=6x subtract 2x from both sides we get -8=4x divide both sides by 4 we get x=-2

OpenStudy (anonymous):

thank you guyz..!